OpenStudy (anonymous):

How does expression written in exponential form: 6^log6(x) turn to a logarithm form: 6^log6(x)=x? HELP!

OpenStudy (anonymous):

M textbook says, 6^log6(x) = y log6(y) = log6(x) y=x 6^log6(x)=y How did they get from 6^log6(x)=y to log6(y) = log6(x)?

OpenStudy (anonymous):

Do you mean \(^6\log(6x)=y\)?

OpenStudy (anonymous):

no, \[6^{\log6x}\]

OpenStudy (anonymous):

And log6(x) is logx with base 6 or log(6x) with base 10?

OpenStudy (anonymous):

logx with base 6

OpenStudy (anonymous):

Then it's the definition of logarithm. if 3^2=9 then log_3(9)=2. 6^log_6(x)=y; say log_6(x)=a 6^a=y then log_6(y)=a log_6(y)=log_6(x)

OpenStudy (anonymous):

In general, a^b=c implies log_a(c)=b